OpenStudy (anonymous):

Speed (S) is measured in meters per second using the formula below, where d is distance traveled in meters and t is time in seconds. S = d/t If an object is moving at a speed of 9 meters per second for a time of 3 seconds, how far did the object travel? A-18 meters B-12 meters C-30 meters D-27 meters

OpenStudy (anonymous):

I dont get what im supose to do. can anybody help me understand this better?

OpenStudy (anonymous):

you just have to substitute the numbers in. So 9 would substitute for s and 3 would substitute for t and then you solve for d

OpenStudy (anonymous):

Im still confused. i get what your sayin but i dont know how to approach the problem.

OpenStudy (anonymous):

what do you mean by that?

OpenStudy (anonymous):

Well we know the speed, and we know the time, so we have to find the distance. Lets solve the formula they gave us for d to isolate the variable: \[S = d/t\] multiply both sides by t\[t(S=d/t)\] and our modified formula looks like this:\[d = st\] What this tells us is that the speed times the time will equal the distance. So 9 * 3 = 27. So D- 27 meters.

OpenStudy (anonymous):

how to figure out D

OpenStudy (anonymous):

ya i guess they are right listen to them

OpenStudy (anonymous):

Does that make sense? You could just enter the values into the formula as is too:\[S=d/t\] \[(9)=d/(3)\]and then solve for d\[27=d\] Whichever way is easiest for you.

OpenStudy (anonymous):

To solve for d, if that part is confusing you, you just have to multiply both sides by 3. Because you have to undo the process being done to the number. So since 3 is being divided (It's in the denominator) you have to multiply by 3 to cancel it out.